OpenStudy (anonymous):

Find the rectangular coordinates of the point with the polar coordinates (8, 3/2Ï€)

OpenStudy (anonymous):

(0, -8) (0, 8) (8, 0) (-8, 0)

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

@Callisto @chihiroasleaf

OpenStudy (chihiroasleaf):

to convert from polar coordinates into Cartesian coordinate Let the coordinates of the point in polar coordinate is \[(r,\theta)\] then \[x = r \cos \theta ; y = r \sin \theta \] from your question, the coordinate is (8, 3/2Ï€) then, what is r? and what is theta? :)